OpenStudy (anonymous):

Q: The tangent to the curve x^3 + x^2y + 4y = 1 at the point (3,-2) has slope...

OpenStudy (raden):

use implicit derivative to derive the curve

OpenStudy (raden):

or u can rewrite it be : x^3 + x^2y + 4y = 1 or x^2y + 4y = 1 - x^3 y(x^2 + 4) = 1 - x^3 y = (1 - x^3)/(x^2+4) use division rule to find dy/dx

OpenStudy (raden):

if y = u/v then dy/dx = (u'v - uv')/(v^2)

OpenStudy (raden):

u = 1-x^3 -----> u' = -3x^2 v = x^2 + 4 -----> v' = 2x just plug of them to the formula above ^ then substitute point x=3 to dy/dx u will get the value of slope

OpenStudy (anonymous):

Thank you for the help!

OpenStudy (raden):

welcome